OpenStudy (curry):

sin(2x)=rad3/2 steps to solve

OpenStudy (curry):

cos4x=-\[\sqrt{3}\]/2

OpenStudy (curry):

-rad3/2=cos4x

OpenStudy (curry):

tan3x=rad3/3

OpenStudy (curry):

sin(2x/3)=-1

OpenStudy (curry):

sin(2x+pi/6)=1/2

OpenStudy (curry):

and tan(x/2)=rad 3

OpenStudy (anonymous):

-rad(3/2)=cos4x=1-2sin^2(2x) \[\sin2x=2\sin^22x-1\] \[2\sin^22x-sinx-1=2\sin^22x-2\sin2x+\sin2x-1=(2\sin2x+1)(\sin2x-1)\] \[\sin2x=1,-1/2\]

OpenStudy (amistre64):

sin(2x)=sqrt.3/2 2x = arcsin(sqrt.3/2) x = arcsin(sqrt.3/2)/2

OpenStudy (curry):

co4x = \[(\sqrt{3})\div2\]

OpenStudy (curry):

i can only ue identitie to solve and i dont think it uses arcsin

OpenStudy (amistre64):

identity for sin and cos is: sin^2+cos^2=1

OpenStudy (curry):

ye i know all the differentys including the double anlge ones the half angle ones and the power reducing ones

OpenStudy (amistre64):

and the identites of: sin(a+b)=sina cosb+sinb cosa cos(a+b) = cosa cosb - sina sinb this tiny little laptop just aint good on for this site :/

OpenStudy (curry):

ye i know those as well

OpenStudy (curry):

u can usew those to derive the double angles cause i did that but how do u solve thee i dont get it cause it syas u can have more 4 anwer for eafch

OpenStudy (amistre64):

the trigs are called circle function since they circle back around time and again to the same "spot" but with a new "name"

OpenStudy (amistre64):

so if you know an angle; the sin and cos parts will also apply for +- multiples of 360 degrees

OpenStudy (amistre64):

drawing rt triangles helps at times too

OpenStudy (curry):

amis i know trig i just dont get how to solve this specific ones, so if u could solve thee then i would get it or atleast the firsst one.

OpenStudy (amistre64):

i cant draw with this thing so ill try the first one at least

OpenStudy (curry):

so for the first one i got pi/3 and pi/6

OpenStudy (amistre64):

sin(2x) = sqrt.3/2 sin = over/hyp ... so 2x = 60 degrees x = 30 degrees as a start

OpenStudy (curry):

but the book syas su can have 4 answers excluding revolutions

OpenStudy (amistre64):

since sin is positive in quads 1 and 2 we have 2x = 60; or 2x = 120

OpenStudy (curry):

5pi/6

OpenStudy (amistre64):

since the periodicity is doubled; we go from a cycle of 2pi,to a cycle of pi

OpenStudy (amistre64):

so we get 4 results; 30 60 210 and 330 i believe plus rotations

OpenStudy (amistre64):

maybe

OpenStudy (amistre64):